Two brass rods AB and BC, each of uniform diameter, will be
brazed together at B to form a nonuniform rod of total length
100 m which will be suspended from a support at A as shown.
Knowing that the density of brass is 8470 kg/m3, determine (a) the
length of rod AB for which the maximum normal stress in ABC is
minimum, (b) the corresponding value of the maximum normal
stress.
Answer #1

A AB= IxQs) 2 176.71mm² = 176.71810-6 m2 ABC= x (100² 78. Symm² 78 suxirt m² b=loo-a WAB - PgAARLAB-8470x 9,81126.71410bxq =
Answer #3

For rod AB

At B we get normal stress due to weight of rod BC

At A we get normal stress due to weight of full rod (AB+BC)

so we can say max stress is at A and min stress at B

then inorder to have min stress we have to consider stress at B should be max normal stress in the rod AB

thats why stress at A=B

Note- Here we are considering min stress at B because it is the weakest point in the rod ABC